How much heat is absorbed by 20 g granite boulder as energy from the sun causes its temperature to change from 10°C to 29°C? (Specific heat capacity of granite is 0.1cal/goC) 38 cal


1
Expert's answer
2022-04-11T18:01:05-0400

Solve for the heat absorbed

Q = m×c×∆T×Q = (20 g)(0.1 cal/g°C)(29 °C - 10 °C) = 38 cal